OpenStudy (anonymous):

Trevor is currently 22 years old. He is half of what his dad's age was six years ago. How old is Trevor's Dad?

OpenStudy (anonymous):

50 years

OpenStudy (anonymous):

T = 22 \[T = \frac{1}{2} (x - 6) \implies 44 = x - 6 \implies x = 50\] Where x = Dad's age.. = 50 years..

OpenStudy (anonymous):

T=22 1/2 dad's before 6 so his dady is 44 before 6 years 44 + 6 =50